# Bilinear spherical maximal function with fractal dilations

### v1

@@ 3. Reduction using Littlewood-Paley decomposition
Let $d\geq 2$. The following pointwise estimates hold.
@@
In view of the lemma above, we are left with proving the desired $L^{p}$-boundedness of the operator $\mathcal{M}^{\mathbf{n}}$, with a suitbale decay in $|\mathbf{n}|$, in order to complete the proof of Theorem 2.1.
@@ 4. Decomposition of multiplier
$\mathcal{A}_{t}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)\widehat{d\sigma}(t\xi,t\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
@@
$\widehat{d\sigma}(t\xi,t\eta)=\frac{2\pi J_{d-1}(|(t\xi,t\eta)|)}{|(t\xi,t\eta)|^{d-1}}.$ (4.1)
@@
$J_{d-1}(r)\sim r^{-\frac{1}{2}}\bigg(e^{2\pi ir}\sum_{j=0}^{\infty}a_{j}r^{-j}+e^{-2\pi ir}\sum_{j=0}^{\infty}a_{j}r^{-j}\bigg)\quad\text{ as }r\to\infty,$
for some coefficients $a_{j}$ and $b_{j}$, such that for all nonnegative integers $N$ and $\alpha$
$\bigg(\frac{d}{dr}\bigg)^{\alpha}\Bigg[J_{d-1}(r)-r^{-\frac{1}{2}}\bigg(e^{2\pi ir}\sum_{j=0}^{N}a_{j}r^{-j}+e^{-2\pi ir}\sum_{j=0}^{N}a_{j}r^{-j}\bigg)\Bigg]=O\big(r^{-\alpha-\frac{N+1}{2}}\big),$ (4.2)
as $r\to\infty$. We refer to [[book-Stein], Chapter VIII] for details.
In view of the asymptotic estimates of Bessel functions, observe that the main term in the multiplier corresponding to the operator $\mathcal{A}_{t}$ behaves like $\frac{e^{2\pi i|(t\xi,t\eta)|}}{|(t\xi,t\eta)|^{\frac{2d-1}{2}}}.$ The remaining terms satisfy higher order decay and hence we need to focus on the main term only. Therefore, we prove the desired $L^{p}$-boundedness of the maximal operator corresponding to bilinear operators defined by
$T^{\mathbf{n}}_{t,k}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{R_{k+n_{1}}f_{1}}(\xi)\widehat{R_{k+n_{2}}f_{2}}(\eta)m(t\xi,t\eta)e^{2\pi it|(\xi,\eta)|}e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
where $t\in E_{k}$ and $m(\xi,\eta)=|(\xi,\eta)|^{-\frac{2d-1}{2}}$.
As described previously, the $L^{p}$-estimates for the global maximal operator are obtained via suitable adaptation of bilinear Calderón-Zygmund theory. We shall restrict ourselves to the case of unit scale, that is $k=0$. The proof is easily adapted to obtain the corresponding estimates for general dyadic scale $2^{k}$.
Let us use the notation $T^{\mathbf{n}}_{t}=T^{\mathbf{n}}_{t,0},$ where $t\in E_{0}$. We fix $t\in E_{0}$ and work with the operator $T^{\mathbf{n}}_{t}.$ Due to the lack of additional decay in $|(\xi,\eta)|$ for the derivative of the term $e^{2\pi it|(\xi,\eta)|},$ we need to decompose this exponential term further.
Let $N$ be a positive integer such that $N\sim 2^{n_{1}}.$ Write the exponential term as follows
$e^{2\pi it|(\xi,\eta)|}=e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi it\sqrt{N^{2}+|\eta|^{2}}},$
@@
Note that $\Phi_{N}$ satisfies the following estimate
$|\partial^{\alpha}_{\eta}\Phi_{N}(\xi,\eta)|\lesssim|\eta|^{-|\alpha|},$ (4.3)
for all multi-indices $\alpha$ uniformly in $\xi$ when $N\leq|\xi|<N+1$.
@@
$F_{N}f(x):=\int_{\mathbb{R}^{d}}\widehat{f}(\xi)\chi_{{}_{N}}(|\xi|)e^{2\pi ix\cdot\xi}d\xi.$
@@
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t}^{\mathbf{n}}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\widehat{F_{N}f_{1}}(\xi)\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
$\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta):=\widehat{f_{2}}(\eta)\widehat{\psi}\left(2^{-n_{2}}\eta\right)e^{2\pi it\left(\sqrt{N^{2}+|\eta|^{2}}\right)},$ (4.4)
@@
$m^{\mathbf{n}}_{t}(\xi,\eta):=m(t\xi,t\eta)\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\zeta}\left(2^{-n_{2}}\eta\right).$
@@
Using the estimate (4.3), we get that
$|\partial^{\alpha}_{\eta}\big(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\big)|\lesssim C_{\alpha}2^{-\frac{2d-1}{2}|\mathbf{n}|}2^{-|\alpha|n_{2}},$ (4.5)
for all multi-indices $\alpha$ uniformly in $t\in E_{0}$.
Since $\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}^{n_{2}}f_{2}(y)e^{-2\pi i\eta\cdot y}dy$, we can rewrite
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{1}}(\xi)U_{t,N}^{n_{2}}f_{2}(y)\Big(\int_{{\mathbb{R}}^{d}}m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\eta\cdot(x-y)}d\eta\Big)e^{2\pi ix\cdot\xi}\;d\xi dy.$ (4.6)
Using integration by parts argument, we obtain
$\int_{{\mathbb{R}}^{d}}m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\eta\cdot(x-y)}d\eta$ (4.7)
$=\frac{1}{(1+4\pi^{2}|2^{n_{2}}(x-y)|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-2^{2n_{2}}\Delta_{\eta})^{\lceil\frac{d}{2}\rceil+1}\Big(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\Big)e^{2\pi i\eta\cdot(x-y)}d\eta,$
@@
$(1-2^{2n_{2}}\Delta_{\eta})e^{2\pi i\eta\cdot(x-y)}=(1+4\pi^{2}|2^{n_{2}}(x-y)|^{2})e^{2\pi i\eta\cdot(x-y)}.$
@@
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{\mathbb{R}^{d}}A_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x,\eta)B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)d\eta,$
@@
$B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta):=$ $\int_{\mathbb{R}^{d}}\left(1-2^{2n_{2}}\Delta_{\eta}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\right)\widehat{F_{N}f_{1}}(\xi)e^{2\pi ix\cdot\xi}d\xi,$ (4.8)
@@
$A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U^{n_{2}}_{t,N}f_{2}(y)e^{2\pi i\eta\cdot(x-y)}}{\left(1+4\pi^{2}\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
@@
$|A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta)|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U^{n_{2}}_{t,N}f_{2}(y)|}{\left(1+4\pi^{2}\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
@@
This gives us the following bound.
$|T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)|\leq L_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta.$ (4.9)
We have the following $L^{p}$-estimates for the terms associated with operators $U_{t,N}^{n_{2}}$ and $B^{\mathbf{n}}_{t,N}$ introduced as above.
@@
Let $a>0$ and $m\in C^{d+2}({\mathbb{R}}^{2d}\setminus\{(0,\dots,0)\}$ be such that
$|\partial_{\xi}^{\alpha}\partial_{\eta}^{\beta}m(\xi,\eta)|\lesssim_{\alpha,\beta}|(\xi,\eta)|^{-a}|\xi|^{-\alpha}|\eta|^{-\beta},$
for all multi-indices $|\alpha|+|\beta|\leq d+2$. For $\mathbf{n}\in{\mathbb{N}}^{2}$ and $N>0$, let $B^{\mathbf{n}}_{s,N}(f)$ be defined as in 4.8. Then, we have
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-|\mathbf{n}|a}2^{n_{2}d}\|f\|_{L^{2}}.$
Lemma 4.2 ([HHY], Lemma 2.2).
Let $I\subset[1,2]$. For $n_{2}\in{\mathbb{N}}$ and $N>0$, let $U_{s,N}^{n_{2}}g$ be defined as in 4.4. Then, we have
@@
The proof of Lemma 4.1 follows by imitating the proof of estimate (2.35) from [HHY] on pages 418-419. Lemma 4.2 was proved for $I=[1,2]$ in [HHY]. The same argument can be adapted for any interval $I\subset[1,2]$ to obtain Lemma 4.2. We skip the details here.
5. Estimate for local maximal operator
We use Bourgain’s interpolation argument to prove restricted weak-type estimates for the local maximal function $\mathcal{M}_{E_{0}}.$ We state the interpolation lemma for convenience. The reader is referred to [Lee1] for more details.
Lemma 5.1 ([Lee1], Lemma 2.6).
Let $\epsilon_{1},\epsilon_{2}>0$. Suppose that $\{T_{j}\}$ is a sequence of linear (or sublinear) operators such that for some $1\leq p_{1},p_{2}<\infty$, and $1\leq q_{1},q_{2}<\infty$,
$\|T_{j}(f)\|_{L^{q_{1}}}\leq M_{1}2^{\epsilon_{1}j}\|f\|_{L^{p_{1}}},~~\|T_{j}(f)\|_{L^{q_{2}}}\leq M_{2}2^{-\epsilon_{2}j}\|f\|_{L^{p_{2}}}.$
Then $T=\sum_{j}T_{j}$ is bounded from $L^{p,1}$ to $L^{q,\infty}$ with
$\|T(f)\|_{L^{q,\infty}}\lesssim M^{\theta}_{1}M^{1-\theta}_{2}\|f\|_{L^{p,1}},$
where $\theta=\frac{\epsilon_{2}}{\epsilon_{1}+\epsilon_{2}}$, $\frac{1}{q}=\frac{\theta}{q_{1}}+\frac{1-\theta}{q_{2}}$ and $\frac{1}{p}=\frac{\theta}{p_{1}}+\frac{1-\theta}{p_{2}}$.

### v2

@@ 3. Reduction using Littlewood-Paley decomposition
Let $d\geq 2$. The following pointwise estimates hold
@@
In view of the lemma above, we are left with proving the desired $L^{p}$-boundedness of the operator $\mathcal{M}^{\mathbf{n}}$, with a suitable decay in $|\mathbf{n}|$, in order to complete the proof of Theorem 2.1.
@@ 4. Decomposition of multiplier
$\mathcal{A}_{t}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)\widehat{d\sigma}(t\xi,t\eta)e^{ix\cdot(\xi+\eta)}\;d\xi d\eta,$
@@
$\widehat{d\sigma}(t\xi,t\eta)=\frac{J_{d-1}(|(t\xi,t\eta)|)}{|(t\xi,t\eta)|^{d-1}}.$ (4.1)
@@
$J_{d-1}(r)\sim r^{-\frac{1}{2}}\bigg(e^{ir}\sum_{j=0}^{\infty}a_{j}r^{-j}+e^{-ir}\sum_{j=0}^{\infty}b_{j}r^{-j}\bigg)\quad\text{ as }r\to\infty,$
for suitable coefficients $a_{j}$ and $b_{j}$, such that for all nonnegative integers $M$ and $\alpha$
$\bigg(\frac{d}{dr}\bigg)^{\alpha}\Bigg[J_{d-1}(r)-r^{-\frac{1}{2}}\bigg(e^{ir}\sum_{j=0}^{M}a_{j}r^{-j}+e^{-ir}\sum_{j=0}^{M}b_{j}r^{-j}\bigg)\Bigg]=O\big(r^{-\alpha-\frac{M+1}{2}}\big),$ (4.2)
as $r\to\infty$. We refer to [[39], page 338] for details.
In view of the asymptotic estimates of Bessel functions, observe that the main term in the multiplier corresponding to the operator $\mathcal{A}_{t}$ behaves like $e^{i|t(\xi,\eta)|}|t(\xi,\eta)|^{-\frac{2d-1}{2}}.$ The remaining terms satisfy higher order decay, and hence we need to focus on the main term only. Therefore, we prove the desired $L^{p}$-boundedness of the maximal operator corresponding to bilinear operators defined by
$T^{\mathbf{n}}_{t,k}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m(t\xi,t\eta)\widehat{\psi}\left(2^{-k-n_{1}}\xi\right)\widehat{\psi}\left(2^{-k-n_{2}}\eta\right)\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}\;d\xi d\eta,$
where $t\in E_{k}$ and $m(\xi,\eta)=e^{it|(\xi,\eta)|}|(\xi,\eta)|^{-\frac{2d-1}{2}}$.
First, we prove the required estimates for the underlying operators for the case of unit scale, that is $k=0$. The analogous results for the general dyadic scale $2^{k}$, are derived using a standard scaling argument.
Let us use the notation $T^{\mathbf{n}}_{t}=T^{\mathbf{n}}_{t,0},$ where $t\in E_{0}$. Due to the lack of additional decay in $|(\xi,\eta)|$ for the derivative of the term $e^{it|(\xi,\eta)|}$ in the multiplier $m$, we need to decompose the exponential term further.
Observe that the support of $\widehat{\psi}$ implies that $2^{n_{1}-1}\leq|\xi|\leq 2^{n_{1}+1}$ and $2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1}$. For a given integer $N\in[2^{n_{1}-1},2^{n_{1}+1}],$ we write the exponential term as follows
$e^{it|(\xi,\eta)|}=e^{it\Phi_{N}(\xi,\eta)}e^{it\sqrt{N^{2}+|\eta|^{2}}},$
@@
Observe that the partial derivative of $\Phi_{N}$ with respect to $\eta$ is given by
$\nabla_{\eta}\Phi_{N}(\xi,\eta)=\frac{-(|\xi|-N)(|\xi|+N)}{\big(\sqrt{|\xi|^{2}+|\eta|^{2}}+\sqrt{N^{2}+|\eta|^{2}}\big)^{2}}\cdot\Big(\frac{\eta}{\sqrt{|\xi|^{2}+|\eta|^{2}}}+\frac{\eta}{\sqrt{N^{2}+|\eta|^{2}}}\Big).$
Therefore, for $\xi$ such that $N\leq|\xi|\leq N+1$, it is easy to verify that the estimate
$|\partial^{\alpha}_{\eta}\Phi_{N}(\xi,\eta)|\lesssim\min\{1,|\eta|^{-|\alpha|}\},$ (4.3)
holds for all multi-indices $\alpha$ uniformly in $\xi$.
@@
$F_{N}f(x):=\int_{\mathbb{R}^{d}}\widehat{f}(\xi){\chi}_{{}_{N}}(|\xi|)e^{ix\cdot\xi}d\xi.$ (4.4)
@@
$S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t,N}^{\mathbf{n}}(\xi,\eta)\widehat{F_{N}f_{1}}(\xi)\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}\;d\xi d\eta,$ (4.5)
$\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta):=\widehat{f_{2}}(\eta)\widehat{\psi}\left(2^{-n_{2}}\eta\right)e^{it\left(\sqrt{N^{2}+|\eta|^{2}}\right)},$ (4.6)
@@
$m_{t,N}^{\mathbf{n}}(\xi,\eta):=\frac{e^{it\Phi_{N}(\xi,\eta)}}{|t(\xi,\eta)|^{\frac{2d-1}{2}}}\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\zeta}\left(2^{-n_{2}}\eta\right).$
@@
Since $\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}^{n_{2}}f_{2}(y)e^{-i\eta\cdot y}dy$, we can rewrite
$S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{1}}(\xi)U_{t,N}^{n_{2}}f_{2}(y)\Big(\int_{{\mathbb{R}}^{d}}m_{t,N}^{\mathbf{n}}(\xi,\eta)e^{i\eta\cdot(x-y)}d\eta\Big)e^{ix\cdot\xi}\;d\xi dy.$ (4.7)
The derivative of $m_{t,N}^{\mathbf{n}}$ is given by
$\partial^{\alpha}_{\eta}\big(m_{t,N}^{\mathbf{n}}(\xi,\eta)\big)=\frac{e^{it\Phi_{N}(\xi,\eta)}}{|t(\xi,\eta)|^{\frac{2d-1}{2}}}\widehat{\psi}\left(2^{-n_{1}}\xi\right)\left(\widehat{\zeta}\left(2^{-n_{2}}\eta\right)\partial^{\alpha}_{\eta}(t\Phi_{N}(\xi,\eta))-\frac{\widehat{\zeta}\left(2^{-n_{2}}\eta\right)}{|(\xi,\eta)|}+2^{-n_{2}}(\partial^{\alpha}\widehat{\zeta})\left(2^{-n_{2}}\eta\right)\right).$
Thus, using the estimate (4.3), we get that
$|\partial^{\alpha}_{\eta}\big(m_{t,N}^{\mathbf{n}}(\xi,\eta)\big)|$ $\lesssim C_{\alpha}2^{-\frac{2d-1}{2}|\mathbf{n}|}2^{-|\alpha|n_{2}},$ (4.8)
for all multi-indices $\alpha$ uniformly in $t\in E_{0}$. Thus, an integration by parts argument gives us
$\int_{{\mathbb{R}}^{d}}m_{t,N}^{\mathbf{n}}(\xi,\eta)e^{i\eta\cdot(x-y)}d\eta$ (4.9)
$=\frac{1}{(1+|2^{n_{2}}(x-y)|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-2^{2n_{2}}\Delta_{\eta})^{\lceil\frac{d}{2}\rceil+1}\Big(m_{t,N}^{\mathbf{n}}(\xi,\eta)\Big)e^{i\eta\cdot(x-y)}d\eta,$
@@
$(1-2^{2n_{2}}\Delta_{\eta})e^{i\eta\cdot(x-y)}=(1+|2^{n_{2}}(x-y)|^{2})e^{i\eta\cdot(x-y)}.$
@@
$S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{\mathbb{R}^{d}}A_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x,\eta)B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)d\eta,$
@@
$B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta):=$ $\int_{\mathbb{R}^{d}}\left(1-2^{2n_{2}}\Delta_{\eta}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m_{t,N}^{\mathbf{n}}(\xi,\eta)\right)\widehat{F_{N}f_{1}}(\xi)e^{ix\cdot\xi}d\xi,$ (4.10)
@@
$A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U^{n_{2}}_{t,N}f_{2}(y)e^{i\eta\cdot(x-y)}}{\left(1+\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
@@
$\left|A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta)\right|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U^{n_{2}}_{t,N}f_{2}(y)|}{\left(1+\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
@@
This gives us the following bound
$|S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)|\leq L_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta.$ (4.11)
The following $L^{p}$-estimates hold for the operators associated with $U_{t,N}^{n_{2}}$ and $B^{\mathbf{n}}_{t,N}$ introduced as above.
@@
Consider the interval $I\subset[1,2]$. For $\mathbf{n}\in{\mathbb{N}}^{2}$ and $N>0$, define the operator $B^{\mathbf{n}}_{s,N}(f)$ as specified in (4.10). Then, we have
$\int_{\mathbb{R}^{d}}\left\|\sup_{s\in I}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|\right\|_{L^{2}}d\eta\lesssim|I|^{\frac{1}{2}}2^{-|\mathbf{n}|\frac{2d-1}{2}}2^{n_{2}d}\|f\|_{L^{2}}.$
Proof.
Using the Sobolev embedding (see [30, Lemma 1.3]), we have
$\left\|\sup_{s\in I}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|\right\|_{L^{2}}$
$\lesssim\left\|\left(\int_{I}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|^{2}ds\right)^{\frac{1}{2}}\right\|_{L^{2}}+\left\|\left(\int_{I}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|^{2}ds\right)^{\frac{1}{2}}\right\|_{L^{2}}^{\frac{1}{2}}\left\|\left(\int_{I}\left|\frac{\partial}{\partial s}B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|^{2}ds\right)^{\frac{1}{2}}\right\|_{L^{2}}^{\frac{1}{2}}.$
Using the estimate (4.8), we have
$\bigg|(1-2^{2n_{2}}\Delta_{\eta})^{\lceil\frac{d}{2}\rceil+1}\Big(m_{s,N}^{\mathbf{n}}(\xi,\eta)\Big)\bigg|$ $\lesssim 2^{-|\mathbf{n}|\frac{2d-1}{2}}{\chi}_{\{2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1}\}}(\eta)$
$\bigg|\frac{\partial}{\partial s}\Big[(1-2^{2n_{2}}\Delta_{\eta})^{\lceil\frac{d}{2}\rceil+1}\big(m_{s,N}^{\mathbf{n}}(\xi,\eta)\big)\Big]\bigg|$ $\lesssim 2^{-|\mathbf{n}|\frac{2d-1}{2}}{\chi}_{\{2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1}\}}(\eta),$
uniformly in $s\in I$ and $\xi$ such that $N\leq\xi<N+1$. Consequently, the Plancherel theorem gives us
$\left\|\left(\int_{I}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|^{2}ds\right)^{\frac{1}{2}}\right\|_{L^{2}}$ $\lesssim|I|^{\frac{1}{2}}2^{-|\mathbf{n}|\frac{2d-1}{2}}{\chi}_{\{2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1}\}}(\eta)\|f\|_{L^{2}}$
$\left\|\left(\int_{I}\left|\frac{\partial}{\partial s}B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|^{2}ds\right)^{\frac{1}{2}}\right\|_{L^{2}}$ $\lesssim|I|^{\frac{1}{2}}2^{-|\mathbf{n}|\frac{2d-1}{2}}{\chi}_{\{2^{n_{2}-1}\leq|\eta|\leq 2^{n_{2}+1}\}}(\eta)\|f\|_{L^{2}}.$
Therefore, by integrating with respect to $\eta$, we complete the proof of Lemma 4.1. ∎
Lemma 4.2 ([24], Lemma 2.2).
Let $I\subset[1,2]$. For $n_{2}\in{\mathbb{N}}$ and $N>0$, let $U_{s,N}^{n_{2}}g$ be defined as in (4.6). Then, we have
@@
Lemma 4.2 was proved for $I=[1,2]$ in [24]. The same argument can be adapted for any interval $I\subset[1,2]$ to obtain the desired bounds in Lemma 4.2. We skip the proof of Lemma 4.2 to avoid repetition. With the decomposition of the multiplier and the results described above, we are in a position to prove the desired $L^{p}$-estimates of the local maximal function $\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})$ in the following section.
5. Estimate for local maximal operator $\mathcal{M}_{E_{0}}$